OpenStudy (anonymous):

A restaurant manager bought 20 packages of bagels. Some packages contained 6 bagels each and the rest contained 12 bagels each. There were 168 bagels in all. How many packages of 12 bagels did the manager buy?

OpenStudy (anonymous):

8 packages of 12 bagels

OpenStudy (anonymous):

\[8 x 12 = 96 \] \[12 x 6 = 72\] \[96 + 72 = 168\]

OpenStudy (jkristia):

one way to solve this is to have 2 variables x and y, x representing box with 6 and y box with 12 then you know that #1; x + y = 20, total number of boxes is 20 #2; 6x + 12y = 168, total number of bagels is 168 now solve for y in #1 so we can replace y in #2 #1; y = 20 - x #2; 6x + 12(20-x) = 168 now solving for x you get 12, so 12 boxes of 6 bagels and #1 y = 20 - 12 = 8 8 boxes of 12 bagels